OpenStudy (anonymous):

2) evaluate each of the following in each case, and explain a. cos^-1(cos(-pi/6) part1: what is the value of cos(-pi/6) part2; to solve the problem cos^-1(cos(-pi/6))find the angle in interval [0,pi] whose cosine is sqrt3/2. b. sin(tan^-1(4/3)) part1: find the measure of the hypotenuse of this triangle. part2: angle (theta) is the angle whose tan is 4/3. find the sine of angle (theta) c. tan^-1(cos(pi/2)) part1: what is the value of cos(pi/2) part2: to solve tan^-1(cos(pi/2)) find the angle in the interval [-pi/2, pi/2] whose tangent equals 0.

OpenStudy (mertsj):

\[\cos (-\frac{\pi}{6})=\frac{\sqrt{3}}{2}\]

OpenStudy (mertsj):

The desired angle is \[\frac{\pi}{6}\]